# Insights of Transitions to Thermoacoustic Instability in Inverse Diffusion Flame using Multifractal Detrended Fluctuation Analysis

### v1

@@ IV.1 Dynamics of IDF for approaching thermoacoustic instability by changing $Re$ (C-1)
During the laminar phase ($Re\sim 1566$), we observe that the acoustic signal ($p^{\prime}$) exhibits low amplitude aperiodic oscillations with a burst of small amplitude for a small time duration (tsimilar-$t\sim$ 7 seconds in Fig. 2 I a). Fast Fourier transformation of the temporal oscillations of the acoustic signal shows multiple significant low-frequency peaks at 100 Hz, 179 Hz, and 264 Hz(Fig. 2 II a). Sen et al. [56] found that the first longitudinal mode of the duct acoustics in the current IDF setup is 100 Hz, determined through an analytical solution following the approach of Dowling [57] and Stow and Dowling [58]. Based on the methodology shown in Sen et al. [29] for determining the non-isothermal and isothermal eigenfrequency modes, we can consider 100 Hz as the isothermal eigenfrequency, while 179 Hz and 264 Hz can be considered as the first and second non-isothermal eigenfrequency modes respectively. On the other hand, the root mean square of acoustic signal (denoted as $p^{\prime}_{rms}$) is observed to be low (as shown at point a of $p^{\prime}_{rms}$ plot in Fig. 2) due to weak flame-acoustic coupling [13]. Besides, the phase-space portrait based on the acoustic fluctuations at $Re\sim 1566$ displays a strange attractor, characterized by complex combinations of phase-space subsets. This suggests that the acoustic fluctuations at such a low $Re$ can exhibit chaotic behavior. Our previous study, using translation error [30], also found that the IDF system demonstrates high-dimensional chaos at low $Re$. We notice a similar dynamics of the system at $Re=2088$.
For shifting to the transitional phase from laminar ($Re\sim 3132$) while the equivalence ratio is kept fixed, we find the combination of low amplitude aperiodic oscillations and the occasional bursts of high amplitude periodic oscillations (see Fig. 2 b I). An earlier study [59] noted similar fluctuations in the acoustic pressure of the multiple flame matrix combustor. Building on this work [59], we classify the fluctuations of the acoustic signal as “mixed mode oscillations”where we can observe one isothermal eigenfrequency mode at 100 Hz and two non-isothermal eigenfrequency modes at 242 Hz and 430 Hz (see Fig. 2 b II). On the other hand, the phase portrait of the acoustic fluctuations exhibits a distinctive structure at the periphery, indicating a significant departure in dynamical features from a chaotic system (see Fig. 2 b III). At $Re\sim 3132$, we notice that $p^{\prime}_{rms}$ increases due to the presence of occasional high amplitude bursts. We continue to notice the mixed mode behavior of the IDF combustor up to $Re=3654$.
@@
During the shifting from the transitional phase to turbulence ($Re=3654$ to $Re=4176$), we observe the appearance of high amplitude periodic oscillations in the acoustic signal (see Fig. 2 c I), signifying the onset of thermoacoustic instability in the IDF combustor. This characteristic continues to be evident during $Re=4176$ to $5533$. $p^{\prime}_{rms}$ becomes high during this period, reflecting greater fluctuations in the acoustic signal and a stronger flame-acoustic coupling in the reacting zone [13]. In Fig. 2 c (I)-(III), we show a representative case of thermoacoustic oscillations corresponding to $Re=5220$. During the fully developed thermoacoustic instability, we find that the non-isothernal eigenfrequency mode dominates at 238 Hz, while the other non-isothermal and isothermal eigenfrequency modes are completely suppressed (see Fig. 2 c (II) ). The phase portrait obtained during thermoacoustic instability clearly exhibits a ring-like structure, signifying the recurrence of the trajectory after a certain time interval. Such a highly deterministic periodic system is observed over a broad range of $Re$ in the turbulent flow regime of the present combustor.
@@ IV.2 Dynamics of IDF for approaching thermoacoustic instability by changing power input (C-2)
Following the MFDFA approach, we vary the structure-function ($F_{q}$) to account for the changes in the measurement scale ($S$) and estimate the slope of $F_{q}$ vs. $S$ for different values of the order of fluctuations ($q$). The resulting slope for different values of $q$, known as the generalized Hurst exponent (noted by $H_{q}$, as mentioned in Sec. III), is discussed in detail in the appendix. As previously mentioned, the slope for $q=2$ is popularly known as the Hurst exponent, which is discussed later. In Fig. 5, we observe the trends of $H_{q}$ (see plots a1 and a2) and mass exponent $\tau_{q}$ (see the plots b1 and b2) that measures the degree of multifractality for different order of fluctuations $q$. In addition, we try to see the multifractal spectrum which illustrates the variation of $f(\alpha)$ with $\alpha$ (see plots a1 and a2 in Fig. 5). From Fig. 5 a1, we notice that $H_{q}$ becomes higher for -5 $\leq q\leq 0$ than for 0 $\leq q\leq 5$ during $Re=1566-2088$ (chaotic regime), indicating that the small-scale oscillations in the system dominate over large-scale fluctuations, a typical characteristic of multifractality. Moreover, the nonlinear relationship of $\tau_{q}$ with $q$ across its range confirms a higher degree of multifractality [60, 61] in the acoustic signal. Consequently, we observe that the singularity spectrum $f(\alpha)$ displays a bell-shaped variation over a wide range of singularity strength $\alpha$.
@@
Figure 7: We perform a recurrence network analysis, as shown in Bhattacharya et al. [30]. Here, we reproduce the variations of the network parameters, namely, global efficiency ($\eta$), degree centrality ($k$) and global clustering coefficient ($C_{G}$) by estimating them on C-1 (source:Chaos 31, 033117 (2021), with a permission from AIP Publishing) [30]. Further, we compare these network measures with $H$ and $\omega$. We show the comparison using dataset C-1 only to avoid the repetition.
Next, we check the variation of $H$ for different dynamical states in Figs. 6 b1 (for the variation in $Re$) and  6 b2 (for the variation in $PI$). We find $H\leq 0.1$ over all the dynamical states in both cases (C-1 and C-2), indicating the long-term switching [62] of the temporal evolutions of the acoustic fluctuations of the system between the higher and lower values. Thus, the trend of the temporal evolutions of the acoustic fluctuations becomes anti-persistent [51] throughout the operating regimes. However, in the current study, our prime interest is to observe the trend of the variation of $H$ at the transitions between the dynamical states. We find a sharp decrease in the measure during the transition from chaotic to mixed mode (C-1 case) and chaotic to weakly ordered state (C-2 case). Further, we find a slow decline in the measure while the system undergoes a transition from mixed mode to thermoacoustic state (C-1 case) and weakly ordered to thermoacoustic state (C-2 case). Similar to $\omega$, the variation of $H$ during the switching between inter-states is found promising. In addition, the gradual decline in $H$ can be suitable to predict the onset of thermoacoustic instability in IDF combustor, as similarly found for the other turbulent combustors [63, 49]. However, note that the change in $H$ from mixed mode to thermoacoustic is noticed to be more prominent than that from weakly ordered to thermoacoustic state. An almost similar trend of $\alpha$ is noticed in Figs. 6 c1 and c2.
In a nutshell, the decreasing trends of $\omega$, $H$ and $\alpha$ prior to the thermoacoustic instability can be useful for identifying the boundaries between the states. However, the threshold for defining these boundaries relies on the control parameters, as the measured values at transitions are not consistent for the two processes. In the next section (Sec. IV.5), we perform a comparative analysis of the proposed measures with existing methods used for IDF in the previous literature [30].
@@
The construction of the recurrence network involves two main steps: first, a phase space reconstruction, and second, the creation of a network based on an adjacency matrix. This adjacency matrix is derived from the recurrence matrix after subtracting the Kronecker delta to eliminate self-loop connections. A more detailed discussion on the network construction is available in the Supplementary Material. We estimate three network parameters, as outlined by our previous investigation [30], which are: global efficiency ($\eta$), degree centrality ($k$) and global clustering coefficient ($C_{G}$). The mathematical formulations for these three measures are provided in the Supplementary Material.
@@ Acknowledgements
S.D. wants to acknowledge the post doctoral fellowship at IIT Madras. S.D. also wants to thank Prof. R I Sujith for his guidance on combustion instabilities. S.D., A.M and S.S. want to thank the Jadavpur University for performing experiments and using lab accessories.

### v3

@@ IV.1 Dynamics of IDF for approaching thermoacoustic instability by changing $Re$ (C-1)
During the laminar phase ($Re\sim 1566$), we observe that the acoustic signal ($p^{\prime}$) exhibits low amplitude aperiodic oscillations with a burst of small amplitude for a small time duration (tsimilar-$t\sim$ 7 seconds in Fig. 2 I a). Fast Fourier transformation of the temporal oscillations of the acoustic signal shows multiple significant low-frequency peaks at 100 Hz, 179 Hz, and 264 Hz(Fig. 2 II a). Sen et al. Sen et al. (2015) found that the first longitudinal mode of the duct acoustics in the current IDF setup is 100 Hz, determined through an analytical solution following the approach of Dowling Dowling (1995) and Stow and Dowling Stow and Dowling (2001). Based on the methodology shown in Sen et al. Sen et al. (2018) for determining the non-isothermal and isothermal eigenfrequency modes, we can consider 100 Hz as the isothermal eigenfrequency, while 179 Hz and 264 Hz can be considered as the first and second non-isothermal eigenfrequency modes respectively. On the other hand, the root mean square of acoustic signal (denoted as $p^{\prime}_{rms}$) is observed to be low (as shown at point a of $p^{\prime}_{rms}$ plot in Fig. 2) due to weak flame-acoustic coupling Sujith and Unni (2020). Besides, the phase-space portrait based on the acoustic fluctuations at $Re\sim 1566$ displays a strange attractor, characterized by complex combinations of phase-space subsets. This suggests that the acoustic fluctuations at such a low $Re$ can exhibit chaotic behavior. Our previous study, using translation error Bhattacharya et al. (2021), also found that the IDF system demonstrates high-dimensional chaos at low $Re$. We notice a similar dynamics of the system at $Re=2088$.
For shifting to the transitional phase from laminar ($Re\sim 3132$) while the equivalence ratio is kept fixed, we find the combination of low amplitude aperiodic oscillations and the occasional bursts of high amplitude periodic oscillations (see Fig. 2 b I). An earlier study Kasthuri et al. (2019b) noted similar fluctuations in the acoustic pressure of the multiple flame matrix combustor. Building on this work Kasthuri et al. (2019b), we classify the fluctuations of the acoustic signal as “mixed mode oscillations”where we can observe one isothermal eigenfrequency mode at 100 Hz and two non-isothermal eigenfrequency modes at 242 Hz and 430 Hz (see Fig. 2 b II). On the other hand, the phase portrait of the acoustic fluctuations exhibits a distinctive structure at the periphery, indicating a significant departure in dynamical features from a chaotic system (see Fig. 2 b III). At $Re\sim 3132$, we notice that $p^{\prime}_{rms}$ increases due to the presence of occasional high amplitude bursts. We continue to notice the mixed mode behavior of the IDF combustor up to $Re=3654$.
@@
During the shifting from the transitional phase to turbulence ($Re=3654$ to $Re=4176$), we observe the appearance of high amplitude periodic oscillations in the acoustic signal (see Fig. 2 c I), signifying the onset of thermoacoustic instability in the IDF combustor. This characteristic continues to be evident during $Re=4176$ to $5533$. $p^{\prime}_{rms}$ becomes high during this period, reflecting greater fluctuations in the acoustic signal and a stronger flame-acoustic coupling in the reacting zone Sujith and Unni (2020); Kasthuri et al. (2022b). In Fig. 2 c (I)-(III), we show a representative case of thermoacoustic oscillations corresponding to $Re=5220$. During the fully developed thermoacoustic instability, we find that the non-isothernal eigenfrequency mode dominates at 238 Hz, while the other non-isothermal and isothermal eigenfrequency modes are completely suppressed (see Fig. 2 c (II) ). The phase portrait obtained during thermoacoustic instability clearly exhibits a ring-like structure, signifying the recurrence of the trajectory after a certain time interval. Such a highly deterministic periodic system is observed over a broad range of $Re$ in the turbulent flow regime of the present combustor.
@@ IV.2 Dynamics of IDF for approaching thermoacoustic instability by changing power input (C-2)
Following the MFDFA approach, we vary the structure-function ($F_{q}$) to account for the changes in the measurement scale ($S$) and estimate the slope of $F_{q}$ vs. $S$ for different values of the order of fluctuations ($q$). The resulting slope for different values of $q$, known as the generalized Hurst exponent (noted by $H_{q}$, as mentioned in Sec. III), is discussed in detail in the appendix. As previously mentioned, the slope for $q=2$ is popularly known as the Hurst exponent, which is discussed later. In Fig. 5, we observe the trends of $H_{q}$ (see plots a1 and a2) and mass exponent $\tau_{q}$ (see the plots b1 and b2) that measures the degree of multifractality for different order of fluctuations $q$. In addition, we try to see the multifractal spectrum which illustrates the variation of $f(\alpha)$ with $\alpha$ (see plots a1 and a2 in Fig. 5). From Fig. 5 a1, we notice that $H_{q}$ becomes higher for -5 $\leq q\leq 0$ than for 0 $\leq q\leq 5$ during $Re=1566-2088$ (chaotic regime), indicating that the small-scale oscillations in the system dominate over large-scale fluctuations, a typical characteristic of multifractality. Moreover, the nonlinear relationship of $\tau_{q}$ with $q$ across its range confirms a higher degree of multifractality Tanna and Pathak (2014); Zhan et al. (2025) in the acoustic signal. Consequently, we observe that the singularity spectrum $f(\alpha)$ displays a bell-shaped variation over a wide range of singularity strength $\alpha$.
@@
Figure 7: We perform a recurrence network analysis, as shown in Bhattacharya et al. Bhattacharya et al. (2021). Here, we reproduce the variations of the network parameters, namely, global efficiency ($\eta$), degree centrality ($k$) and global clustering coefficient ($C_{G}$) by estimating them on C-1 (source:Chaos 31, 033117 (2021), with a permission from AIP Publishing) Bhattacharya et al. (2021). Further, we compare these network measures with $H$ and $\omega$. We show the comparison using dataset C-1 only to avoid the repetition.
Next, we check the variation of $H$ for different dynamical states in Figs. 6 b1 (for the variation in $Re$) and  6 b2 (for the variation in $PI$). We find $H\leq 0.1$ over all the dynamical states in both cases (C-1 and C-2), indicating the long-term switching Alvarez-Ramirez et al. (2008) of the temporal evolutions of the acoustic fluctuations of the system between the higher and lower values. Thus, the trend of the temporal evolutions of the acoustic fluctuations becomes anti-persistent Ihlen (2012) throughout the operating regimes. However, in the current study, our prime interest is to observe the trend of the variation of $H$ at the transitions between the dynamical states. We find a sharp decrease in the measure during the transition from chaotic to mixed mode (C-1 case) and chaotic to weakly ordered state (C-2 case). Further, we find a slow decline in the measure while the system undergoes a transition from mixed mode to thermoacoustic state (C-1 case) and weakly ordered to thermoacoustic state (C-2 case). Similar to $\omega$, the variation of $H$ during the switching between inter-states is found promising. In addition, the gradual decline in $H$ can be suitable to predict the onset of thermoacoustic instability in IDF combustor, as similarly found for the other turbulent combustors Pavithran et al. (2020); Nair and Sujith (2014). However, note that the change in $H$ from mixed mode to thermoacoustic is noticed to be more prominent than that from weakly ordered to thermoacoustic state. An almost similar trend of $\alpha$ is noticed in Figs. 6 c1 and c2.
In a nutshell, the decreasing trends of $\omega$, $H$ and $\alpha$ prior to the thermoacoustic instability can be useful for identifying the boundaries between the states. However, the threshold for defining these boundaries relies on the control parameters, as the measured values at transitions are not consistent for the two processes. In the next section (Sec. IV.5), we perform a comparative analysis of the proposed measures with existing methods used for IDF in the previous literature Bhattacharya et al. (2021).
@@
The construction of the recurrence network involves two main steps: first, a phase space reconstruction, and second, the creation of a network based on an adjacency matrix. This adjacency matrix is derived from the recurrence matrix after subtracting the Kronecker delta to eliminate self-loop connections. A more detailed discussion on the network construction is available in the Supplementary Material. We estimate three network parameters, as outlined by our previous investigation Bhattacharya et al. (2021), which are: global efficiency ($\eta$), degree centrality ($k$) and global clustering coefficient ($C_{G}$). The mathematical formulations for these three measures are provided in the Supplementary Material.
@@ V Conclusions
Supplementary Material
We include Supplementary Material that clearly outlines the methods of phase reconstruction, recurrence networks, and the 0-1 test. The performance of the 0-1 test for the C-2 case is also detailed in the Supplementary Material.
@@ Acknowledgements
S.D. wants to thank Prof. R I Sujith for his guidance on combustion instabilities. S.D. also wants to thank the combustion lab at Jadavpur University for performing experiments and using lab accessories.
Data Availability Statement
The data that support the findings of this study are available from the corresponding author upon reasonable request.
Declarations: Conflict of interests
The authors declare no competing interests.